The space shuttle releases a satellite into a circular orbit 630 km

km above the Earth.

How fast must the shuttle be moving (relative to Earth's center) when the release occurs?


1
Expert's answer
2021-12-12T16:41:01-0500
"v=\sqrt{(6.67\cdot10^{-11})\frac{(5.97\cdot10^{24})}{((630+6371)\cdot10^{3})}}\\v=7542\frac{m}{s}"